# Fearless Stochasticity in Expectation Propagation

### v1

@@ Abstract
Expectation propagation (EP) is a family of algorithms for performing approximate inference in probabilistic models. The updates of EP involve the evaluation of moments—expectations of certain functions—which can be estimated from Monte Carlo (MC) samples. However, the updates are not robust to MC noise when performed naively, and various prior works have attempted to address this issue in different ways. In this work, we provide a novel perspective on the moment-matching updates of EP; namely, that they perform natural-gradient-based optimisation of a variational objective. We use this insight to motivate two new EP variants, with updates that are particularly well-suited to MC estimation; they remain stable and are most sample-efficient when estimated with just a single sample. These new variants combine the benefits of their predecessors and address key weaknesses. In particular, they are easier to tune, offer an improved speed-accuracy trade-off, and do not rely on the use of debiasing estimators. We demonstrate their efficacy on a variety of probabilistic inference tasks.
@@ 1 Introduction
EP involves the evaluation of moments—expectations of certain functions—under distributions that are derived from the model of interest. EP is usually applied to models for which these moments have convenient closed-form expressions or can be accurately estimated using deterministic methods. Moments can also be estimated using Monte Carlo (MC) samples, significantly expanding the set of models EP can be applied to; however, the updates of EP are not robust to MC noise when performed naively, and various prior works have attempted to address this issue in different ways [14, 44, 47].
@@ 2 Background
Let $\mathcal{F}$ be the tractable, minimal exponential family of distributions (see Appendix A), defined by the statistic function $s(.)$ with respect to base measure $\nu(.)$. Let $\Omega$, $\mathcal{M}$, $A(.)$ and $A^{*}(.)$ denote the natural domain, mean domain, log partition function, and negative entropy function of $\mathcal{F}$, respectively.
Let $p_{0}$ be the member of $\mathcal{F}$ with natural parameter $\theta_{0}$, so that $p_{0}(z)=\exp\bm{(}\theta_{0}^{\top}s(z)-A(\theta_{0})\bm{)}$ [^1] We use e.g. $p$ to refer to a distribution, and $p(.)$ or $p(z)$ for its density, throughout., and assume that a distribution of interest $p^{*}$, the target distribution, has a density of the form
@@
In Bayesian inference settings, $p^{*}$ would be the posterior distribution over parameters $z$ given some observed data $\mathcal{D}$, where $p_{0}$ may correspond to a prior distribution, and $\{\ell_{i}(z)\}_{i}$ to log-likelihood terms given some partition of $\mathcal{D}$ [^2] Going forward, we assume $i$ and $j$ to be taken from the index set $\{1,\ldots,m\}$ unless otherwise specified. However, we consider the more general setting in which the target distribution is simply a product of factors; note that we can assume form (1) without loss of generality if we allow $p_{0}$ to be improper. The inference problem typically amounts to evaluating quantities derived from the normalised density $p^{*}(z)$, such as samples, summary statistics, or expectations of given functions. In some special cases, these quantities can be computed exactly, but this is not feasible in general, and approximations must be employed.
@@ 2.1 Expectation propagation (EP)
By assumption $\mathcal{F}$ is tractable, and so provided that $(\theta_{0}+\sum_{i=1}^{m}\lambda_{i})\in\Omega$, $p$ is a tractable member of $\mathcal{F}$. Each factor $\exp\bm{(}\lambda_{i}^{\top}s(z)\bm{)}$ is known as a site potential, and can be roughly interpreted as a $\mathcal{F}$-approximation to the $i$-th target factor, $\exp\bm{(}\ell_{i}(z)\bm{)}$ [^3] This interpretation is not precise, since it is not necessarily the case that $\lambda_{i}\in\Omega$. $\lambda_{i}$ is known as the $i$-th site parameter.
@@
Let the $i$-th locally extended family, denoted $\mathcal{F}_{i}$, be the exponential family defined by the statistic function $s_{i}(z)=(s(z),\ell_{i}(z))$ with respect to base measure $\nu(.)$. Let $\Theta_{i},\mathcal{M}_{i}$ and $A_{i}(.)$ denote the natural domain, mean domain, and log partition function of $\mathcal{F}_{i}$, respectively. Note that a member of $\mathcal{F}_{i}$ roughly corresponds to a distribution whose density is the (normalised) product of a member of $\mathcal{F}$ with the $i$-th target factor raised to a power. Unlike $\mathcal{F}$, we do not assume $\mathcal{F}_{i}$ is minimal.
@@
$\max_{\{\theta_{i}\in\Omega\}_{i}}\min_{\{\lambda_{i}\}_{i}}L(\theta_{1},\ldots,\theta_{m},\lambda_{1},\ldots,\lambda_{m}),\text{ where }$
$L(\theta_{1},\ldots,\theta_{m},\lambda_{1},\ldots,\lambda_{m})=A\Big{(}\theta_{0}+\sum_{i}\lambda_{i}\Big{)}+\sum_{i}\beta_{i}\big{[}A_{i}\bm{(}(\theta_{i}-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}-A(\theta_{i})\big{]}.$ (3)
@@
EP [33, 36], power EP [30] and double-loop EP [21], can all be viewed as alternating between some number ($\geq 1$) of inner updates to decrease $L$ with respect to $\{\lambda_{i}\}_{i}$, with an outer update to increase $L$ with respect to $\{\theta_{i}\}_{i}$ [14, 22]. EP is not typically presented in this way, but by doing so we will be able to present a unified algorithm that succinctly illustrates the relationship between the different variants. The inner and outer updates are given by
Inner update: $\lambda_{i}\leftarrow\lambda_{i}-\alpha\Big{(}\theta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\bm{(}\mathbb{E}_{p_{i}(z)}[s(z)]\bm{)}\Big{)},$ (4)
Outer update: $\theta_{i}\leftarrow\theta_{0}+\sum_{j}\lambda_{j},$ (5)
@@
$p_{i}(z)$ $\propto p_{0}(z)\exp\Big((\theta_{i}-\beta^{-1}\lambda_{i})^{\top}s(z)+\beta^{-1}\ell_{i}(z)\Big{missing}).$ (6)
The hyperparameter $\alpha$ controls the level of damping, which is used to aid convergence. An undamped inner update (with $\alpha=1$) can be seen as performing moment matching between $p$ and $p_{i}$. The inner updates can be performed either serially (over $i$) or in parallel. In this work we assume they are applied in parallel, but the ideas presented can easily be extended to the serial case. Heskes and Zoeter [22] showed that (4) follows a decrease direction in $L$ with respect to $\lambda_{i}$, and so is guaranteed to decrease $L$ when $\alpha$ is small enough. See Appendix B for a derivation of the inner update. When $\{\lambda_{i}\}_{i}$ are at a partial minimum of $L$ (for fixed $\{\theta_{i}\}_{i}$), the outer update performs an exact partial maximisation of the primal form of the variational problem; see Hasenclever et al. [14] for details.
@@
The double-loop EP algorithm of Heskes and Zoeter [21] repeats the inner update to convergence before performing each outer update, which ensures convergence of the overall procedure. The usual presentation of EP combines (4) and (5) into a single update in $\lambda_{i}$; however, Jylänki et al. [27] observed that it can also be viewed as performing double-loop EP with just a single inner step per outer update (which is not guaranteed to converge in general). By taking this viewpoint, we are we are able to present EP, power EP, and their double-loop counterparts as a single algorithm, presented in Algorithm 1. Note that the dependence on $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta_{i}}\}_{i}$ comes through the definition of $p_{i}(.)$.
Algorithm 1 EP ($\beta_{i}$=$1$,$n_{\text{inner}}$=$1$), power EP ($\beta_{i}\neq 1$,$n_{\text{inner}}$=$1$), and their double-loop variants ($n_{\text{inner}}$>$1$)
$\mathcal{F}$, $\theta_{0}$, $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta_{i}}\}_{i}$, $\{\ell_{i}(z)\}_{i}$, $\{\lambda_{i}\}_{i}$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{\text{inner}}}$, $\alpha$
@@
     $\theta_{i}\leftarrow\theta_{0}+\sum_{j}\lambda_{j}$ for $i=1$ to $m$
@@
         for $i=1$ to $m$ in parallel do
              $\lambda_{i}\leftarrow\lambda_{i}-\alpha\Big{(}\theta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\big{(}\mathbb{E}_{p_{i}(z)}[s(z)]\big{)}\Big{)}$
@@
We can think of "exact" double-loop EP as corresponding to ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{\text{inner}}}=\infty$, with the inner loop exiting only once some convergence criteria has been satisfied. In practice, a truncated inner loop is often used by setting ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{\text{inner}}}$ to some small number [14, 27]. Upon convergence of Algorithm 1, the approximation $p(z)\approx p^{*}(z)$ is given by (2). Going forwards we will refer to the family of algorithms encompassed by Algorithm 1 simply as EP.
@@
EP is typically applied to models for which the updates either have closed-form expressions, or can be accurately estimated using deterministic numerical methods. However, as update (4) only depends on the target distribution through expectations under the tilted distributions, this suggests that updates could also be performed using sampled estimates of those expectations. By using MC methods to estimate the tilted distribution moments, EP can be applied in a black-box manner, dramatically expanding the space of models it can be applied to. Instead of performing a single large sampling task—as would be required by applying MC methods directly—EP can instead solve several simpler ones, gaining significant computational advantages [14, 44, 47]. Unfortunately, when performed naively, update (4) is not robust to MC noise. This is because the moment estimates are converted to the natural (site) parameter space by mapping through $\nabla A^{*}(.)$, which is not linear in general, and so MC noise in the estimates leads to biased updates of $\lambda_{i}$.
@@ 3.1 Natural gradient view of EP
Consider instead $\tilde{p}_{i}^{(t)}\in\mathcal{F}$ with natural parameter $\eta_{i}^{(t)}(\lambda_{i})=\theta_{0}+\sum_{j}\lambda_{j}^{(t)}+\alpha^{-1}(\lambda_{i}-\lambda_{i}^{(t)})$. The $t$ superscript on the site parameters indicates that they are fixed for the current iteration, and so the distribution is fully parameterised by $\lambda_{i}$. Note that when $\lambda_{j}=\lambda_{j}^{(t)}\ \forall\ j$, we have $\tilde{p}_{i}^{(t)}=p$. Proposition 1, which we prove in Appendix C, states that the moment-matching updates of EP can be viewed as performing NGD in $L$ with respect to mean parameters of $\tilde{p}_{i}^{(t)}$.
@@
For $\alpha>0$, the moment-matching update of EP (4) is equivalent to performing an NGD step in $L$ with respect to the mean parameters of $\tilde{p}_{i}^{(t)}$ with step size $\alpha^{-1}$. That is, for $\mu_{i}=\mathbb{E}_{\tilde{p}_{i}^{(t)}(z)}[s(z)]$, and $\tilde{F}^{(t)}_{i}(\mu_{i})$ the FIM of $\tilde{p}_{i}^{(t)}$ with respect to $\mu_{i}$, we have
$\mu_{i}-\alpha^{-1}\Big{[}\tilde{F}_{i}^{(t)}(\mu_{i})\Big{]}^{-1}\partialderivative{L}{\mu_{i}}$ $=\Big{(}\nabla A\circ\eta_{i}^{(t)}\Big{)}\bigg{(}\lambda_{i}-\alpha\Big{(}\theta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\bm{(}\mathbb{E}_{p_{i}(z)}[s(z)]\bm{)}\Big{)}\bigg{)}.$ (7)
Note that the right hand side of (7) just maps update (4) to mean parameters of $\tilde{p}_{i}^{(t)}$. [^4] Note the apparent contradiction that by decreasing $\alpha$ (increasing the damping) we actually increase the NGD step size. This is resolved by observing that the definition of $\tilde{p}^{(t)}_{i}$ also changes with $\alpha$. We discussed in Section 2 that noise in the moment estimates results in bias in the update to $\lambda_{i}$ of EP, due to the noise being passed through the nonlinear map $\nabla A^{*}(.)$. This alone would not be a problem if the updates followed unbiased descent direction estimates in some other fixed parameterisation. Let $\eta_{i}=\nabla A^{*}(\mu_{i})$; then, from the definition of $L$ and $\mu_{i}$, we have
$\partialderivative{L}{\mu_{i}}$ $=\partialderivative{\eta_{i}}{\mu_{i}}\partialderivative{\lambda_{i}}{\eta_{i}}\Big{(}\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)},$ (8)
and so when $\mathbb{E}_{p_{i}(z)}[s(z)]$ is estimated with unbiased samples, the NGD update (7) in $\mu_{i}$ is also unbiased. However, a sequence of updates introduces bias. This is because the parameterisation changes from one update to the next, with the mean parameters of $\tilde{p}_{i}^{(t)}$ mapped to those of $\tilde{p}_{i}^{(t+1)}$ by
@@
where $\big{(}\eta_{i}^{(t)}\big{)}^{-1}(.)$ is the inverse of $\eta_{i}^{(t)}(.)$, and $\eta_{i}^{(t+1)}(.)$ is defined using the site parameters after the update at time $t$. This map is not linear in general, and so an unbiased, noisy update of $\mu_{i}$ at one time step, results in bias when mapped to the next. This bias makes the EP updates unstable in the presence of MC noise. This has led previous work to use methods, specific to $\mathcal{F}$, for obtaining approximately unbiased natural parameter estimates from samples. However, even with these debiasing methods, a relatively large number of samples is still needed in practice [44, 47].
We now use the natural gradient interpretation of the EP updates to motivate two new variants that are far more robust to MC noise; in particular, they are stable, and most sample-efficient, when updates are estimated with just a single sample. Furthermore, unlike methods that rely on deibasing estimators, they do not require sample thinning (see Section 3.4). This largely eliminates two hyperparameters for the practitioner. While practical considerations, such as ensuring efficient use of parallel hardware, may justify using more than one sample per update, this can in principle be tuned a priori, much like the batch size hyperparameter of stochastic gradient descent.

### v2

@@ Abstract
Expectation propagation (EP) is a family of algorithms for performing approximate inference in probabilistic models. The updates of EP involve the evaluation of moments—expectations of certain functions—which can be estimated from Monte Carlo (MC) samples. However, the updates are not robust to MC noise when performed naively, and various prior works have attempted to address this issue in different ways. In this work, we provide a novel perspective on the moment-matching updates of EP; namely, that they perform natural-gradient-based optimisation of a variational objective. We use this insight to motivate two new EP variants, with updates that are particularly well-suited to MC estimation. They remain stable and are most sample-efficient when estimated with just a single sample. These new variants combine the benefits of their predecessors and address key weaknesses. In particular, they are easier to tune, offer an improved speed-accuracy trade-off, and do not rely on the use of debiasing estimators. We demonstrate their efficacy on a variety of probabilistic inference tasks.
@@ 1 Introduction
EP involves the evaluation of moments—expectations of certain functions—under distributions that are derived from the model of interest. EP is usually applied to models for which these moments have convenient closed-form expressions or can be accurately estimated using deterministic methods. Moments can also be estimated using Monte Carlo (MC) samples, significantly expanding the set of models EP can be applied to. However, the updates of EP are not robust to MC noise when performed naively, and various prior works have attempted to address this issue in different ways [17, 48, 52].
@@ 2 Background
Let $\mathcal{F}$ be the tractable, minimal exponential family of distributions (see Appendix A), defined by the statistic function $s(.)$ with respect to base measure $\nu(.)$. Let $\Omega$, $\mathcal{M}$, $A(.)$ denote the natural domain, mean domain and log-partition function of $\mathcal{F}$, respectively, and $A^{*}(.)$ the convex dual of $A(.)$.
Let $p_{0}$ be the member of $\mathcal{F}$ with natural parameter $\eta_{0}$, so that $p_{0}(z)=\exp\bm{(}\eta_{0}^{\top}s(z)-A(\eta_{0})\bm{)}$ [^1] We use e.g. $p$ to refer to a distribution, and $p(.)$ or $p(z)$ for its density, throughout., and assume that a distribution of interest $p^{*}$, the target distribution, has a density of the form
@@
In Bayesian inference settings, $p^{*}$ would be the posterior distribution over parameters $z$ given some observed data $\mathcal{D}$, where $p_{0}$ may correspond to a prior distribution, and $\{\ell_{i}(z)\}_{i}$ to log-likelihood terms given some partition of $\mathcal{D}$ [^2] Going forward, we assume $i$ and $j$ to be taken from the index set $\{1,\ldots,m\}$ unless otherwise specified. However, we consider the more general setting in which the target distribution is simply a product of factors. Note that we can assume form (1) without loss of generality if we allow $p_{0}$ to be improper. The inference problem typically amounts to computing quantities derived from the normalised density $p^{*}(z)$, such as samples, summary statistics, or expectations of given functions. In some special cases, these quantities can be computed exactly, but this is not feasible in general, and approximations must be employed.
@@ 2.1 Expectation propagation (EP)
By assumption $\mathcal{F}$ is tractable, and so provided that $\smash{(\eta_{0}+\sum_{i}\lambda_{i})\in\Omega}$, $p$ is a tractable member of $\mathcal{F}$. Each factor $\smash{\exp\bm{(}\lambda_{i}^{\top}s(z)\bm{)}}$ is known as a site potential, and can be roughly interpreted as a $\mathcal{F}$-approximation to the $i$-th target factor, $\exp\bm{(}\ell_{i}(z)\bm{)}$ [^3] This interpretation is not precise, since it is not necessarily the case that $\lambda_{i}\in\Omega$. $\lambda_{i}$ is known as the $i$-th site parameter.
@@
Let the $i$-th locally extended family, denoted $\mathcal{F}_{i}$, be the exponential family defined by the statistic function $s_{i}(z)=(s(z),\ell_{i}(z))$ with respect to base measure $\nu(.)$. Let $\Omega_{i},\mathcal{M}_{i}$ and $A_{i}(.)$ denote the natural domain, mean domain, and log partition function of $\mathcal{F}_{i}$, respectively. Note that a member of $\mathcal{F}_{i}$ roughly corresponds to a distribution whose density is the (normalised) product of a member of $\mathcal{F}$ with the $i$-th target factor raised to a power. Unlike $\mathcal{F}$, we do not assume $\mathcal{F}_{i}$ is minimal.
@@
$\max_{\theta\in\Omega}\min_{\{\lambda_{i}\}_{i}}L(\theta,\lambda_{1},\ldots,\lambda_{m}),\text{ where }$
$L(\theta,\lambda_{1},\ldots,\lambda_{m})=A\Big{(}\eta_{0}+\sum_{i}\lambda_{i}\Big{)}+\sum_{i}\beta_{i}\big{[}A_{i}\bm{(}(\theta-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}-A(\theta)\big{]}.$ (3)
@@
EP [37, 40], power EP [34] and double-loop EP [24], can all be viewed as alternating between some number ($\geq 1$) of inner updates to decrease $L$ with respect to $\{\lambda_{i}\}_{i}$, with an outer update to increase $L$ with respect to $\theta$ [17, 25]. EP is not typically presented in this way, but by doing so we will be able to present a unified algorithm that succinctly illustrates the relationship between the different variants. We show equivalence with the conventional presentation of EP in Appendix B. The inner and outer updates are given by
Inner update: $\lambda_{i}\leftarrow\lambda_{i}-\alpha\Big{(}\eta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\bm{(}\mathbb{E}_{p_{i}(z)}[s(z)]\bm{)}\Big{)},$ (4)
Outer update: $\theta\leftarrow\eta_{0}+\sum_{j}\lambda_{j},$ (5)
@@
$p_{i}(z)$ $\propto\exp\left((\theta-\beta_{i}^{-1}\lambda_{i})^{\top}s(z)+\beta_{i}^{-1}\ell_{i}(z)\right).$ (6)
The hyperparameter $\alpha$ controls the level of damping, which is used to aid convergence. An undamped inner update (with $\alpha=1$) can be seen as performing moment matching between $p$ and $p_{i}$. The expectation in (4) is most often computed analytically, or estimated using deterministic numerical methods. It can also be estimated by sampling from $p_{i}$, however, we will later show that the resulting stochasticity can lead to a biased and unstable procedure. The inner updates can be performed either serially or in parallel (over $i$). In this work we assume they are applied in parallel, but the ideas presented can easily be extended to the serial case. Heskes and Zoeter [25] showed that (4) follows a decrease direction in $L$ with respect to $\lambda_{i}$, and so is guaranteed to decrease $L$ when $\alpha$ is small enough. See Appendix C for a derivation of the inner update. When $\{\lambda_{i}\}_{i}$ are at a partial minimum of $L$ (for fixed $\theta$), the outer update performs an exact partial maximisation of the primal form of the variational problem – see Hasenclever et al. [17] for details. All of the EP variants considered in this paper differ only in their handling of the inner minimisation, and this is our focus.
@@
The double-loop EP algorithm of Heskes and Zoeter [24] repeats the inner update to convergence before performing each outer update, which ensures convergence of the overall procedure. The usual presentation of EP combines (4) and (5) into a single update in $\lambda_{i}$ (see Appendix B), however, Jylänki et al. [30] observed that it can also be viewed as performing double-loop EP with just a single inner update per outer update (which is not guaranteed to converge in general). By taking this view, we are we are able to present EP, power EP, and their double-loop counterparts as a single algorithm, presented in Algorithm 1. We do so primarily to illustrate how these variants are related to one another, and to the new variants of Section 3.
Algorithm 1 EP ($\beta_{i}$=$1$,$n_{\text{inner}}$=$1$), power EP ($\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\beta_{i}{\neq}1$,$n_{\text{inner}}$=$1$), and their double-loop variants ($n_{\text{inner}}$>$1$)
$\mathcal{F}$, $\eta_{0}$, $\{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\beta_{i}}\}_{i}$, $\{\ell_{i}(z)\}_{i}$, $\{\lambda_{i}\}_{i}$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{\text{inner}}}$, $\alpha$
@@
     $\smash{\theta\leftarrow\eta_{0}+\cramped{\sum}_{j}\lambda_{j}}$
@@
         for $i=1$ to $m$ in parallel do Stochastic estimation of this expectation can lead to biased and unstable updates.
              $\smash{\lambda_{i}\leftarrow\lambda_{i}-\alpha(\eta_{0}+\cramped{\sum}_{j}\lambda_{j}-\nabla A^{*}\bm{(}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{p_{i}(z)}[s(z)]}\bm{)})}$
@@
Note that dependence on $\{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\beta_{i}}\}_{i}$ comes through the definition of $p_{i}(.)$. We can think of “exact” double-loop EP as corresponding to ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{\text{inner}}=\infty}$, with the inner loop exiting once some convergence criterion has been satisfied. In practice, a truncated inner loop is often used by setting ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{\text{inner}}}$ to some small number [17, 30]. Upon convergence of Algorithm 1, the approximation $p(z)\approx p^{*}(z)$ is given by (2). Going forwards we will refer to the family of algorithms encompassed by Algorithm 1 simply as EP.
@@
EP is typically applied to models for which the updates either have closed-form expressions, or can be accurately estimated using deterministic numerical methods. However, as update (4) only depends on the target distribution through expectations under the tilted distributions, this suggests that updates could also be performed using sampled estimates of those expectations. By using MC methods to estimate the tilted distribution moments, EP can be used in a black-box manner, dramatically expanding the set of models it can be applied to. Instead of performing a single large sampling task—as would be required by applying MC methods directly—EP can instead solve several simpler ones, gaining significant computational advantages [17, 48, 52]. Unfortunately, when performed naively, update (4) is not robust to MC noise. This is because the moment estimates are converted to the natural (site) parameter space by mapping through $\nabla A^{*}(.)$, which is not linear in general, and so MC noise in the estimates leads to biased updates of $\lambda_{i}$.
@@ 3.1 Natural gradient view of EP
Consider instead $\cramped{\tilde{p}_{i}}^{(t)}\in\mathcal{F}$ with natural parameter $\smash{\cramped{\eta_{i}}^{(t)}(\lambda_{i})=\eta_{0}+\sum_{j}\cramped{\lambda_{j}}^{(t)}+\alpha^{-1}(\lambda_{i}-\cramped{\lambda_{i}}^{(t)})}$. The $t$ superscript on the site parameters indicates that they are fixed for the current iteration, and so the distribution is fully parameterised by $\lambda_{i}$. Note that when $\smash{\lambda_{j}=\cramped{\lambda_{j}}^{(t)}\ \forall\ j}$, we have $\smash{\cramped{\tilde{p}_{i}}^{(t)}=p}$. Proposition 1, which we prove in Appendix D, states that the moment-matching updates of EP can be viewed as performing NGD in $L$ with respect to mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$.
@@
For $\alpha>0$, the moment-matching update of EP (4) is equivalent to performing an NGD step in $L$ with respect to the mean parameters of $\cramped{\tilde{p}_{i}}^{(t)}$ with step size $\alpha^{-1}$. That is, for $\mu_{i}=\mathbb{E}_{\cramped{\tilde{p}_{i}}^{(t)}(z)}[s(z)]$, and $\smash{\cramped{\tilde{F}_{i}}^{(t)}(\mu_{i})}$ the FIM of $\cramped{\tilde{p}_{i}}^{(t)}$ with respect to $\mu_{i}$, we have
$\mu_{i}-\alpha^{-1}\Big{[}\tilde{F}_{i}^{(t)}(\mu_{i})\Big{]}^{-1}\partialderivative{L}{\mu_{i}}$ $=\Big{(}\nabla A\circ\eta_{i}^{(t)}\Big{)}\bigg{(}\lambda_{i}-\alpha\Big{(}\eta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\bm{(}\mathbb{E}_{p_{i}(z)}[s(z)]\bm{)}\Big{)}\bigg{)}.$ (7)
Note that the right hand side of (7) just maps update (4) to mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$.  [^4] Note the apparent contradiction that by decreasing $\alpha$ (increasing the damping) we actually increase the NGD step size. This is resolved by observing that the definition of $\tilde{p}^{(t)}_{i}$ also changes with $\alpha$. We discussed in Section 2 that noise in the moment estimates results in bias in the update to $\lambda_{i}$ of EP, due to the noise being passed through the nonlinear map $\nabla A^{*}(.)$. This alone would not be a problem if the updates followed unbiased descent direction estimates in some other fixed parameterisation. Let $\eta_{i}=\nabla A^{*}(\mu_{i})$, then, from the definition of $L$ and $\mu_{i}$, we have
$\partialderivative{L}{\mu_{i}}$ $=\partialderivative{\eta_{i}}{\mu_{i}}\partialderivative{\lambda_{i}}{\eta_{i}}\Big{(}\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)},$ (8)
and so when $\mathbb{E}_{p_{i}(z)}[s(z)]$ is estimated with unbiased samples, the NGD update (7) in $\mu_{i}$ is also unbiased. However, a sequence of updates introduces bias. This is because the parameterisation changes from one update to the next, with the mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ mapped to those of $\smash{\cramped{\tilde{p}_{i}}^{(t+1)}}$ by
@@
where $\smash{\cramped{(\cramped{\eta_{i}}^{(t)})}^{-1}(.)}$ is the inverse of $\smash{\cramped{\eta_{i}}^{(t)}(.)}$, and $\smash{\cramped{\eta_{i}}^{(t+1)}(.)}$ is defined using the site parameters after the update at time $t$. This map is not linear in general, and so an unbiased, noisy update of $\mu_{i}$ at one time step, results in bias when mapped to the next. This bias can make the EP updates unstable in the presence of MC noise, leading previous work to use methods, specific to $\mathcal{F}$, for obtaining approximately unbiased natural parameter estimates from samples. However, even with these debiasing methods, a relatively large number of samples is still needed in practice [48, 52]. Here we have assumed parallel application of (4), but the bias at site $i$ is induced whenever the parameters for other sites change from one update to the next, which occurs in both parallel and serial settings.
We now use the natural gradient interpretation of the EP updates to motivate two new variants that are far more robust to MC noise. In particular, they are stable, and most sample-efficient, when updates are estimated with just a single sample. Furthermore, unlike methods that rely on deibasing estimators, they do not require sample thinning (see Section 3.4). This largely eliminates two hyperparameters for the practitioner. While practical considerations, such as ensuring efficient use of parallel hardware, may justify using more than one sample per update, this can in principle be tuned a priori, much like the batch size hyperparameter of stochastic gradient descent.